18-3.8t=7036-1.9t

i need to find the answer to T.thank you

Answer by LilSkittleMd(119)

18-3.8t=7036-1.9t
First you need to get the variables on one side and the numbers with no variables on the other side. Subtract 18 from both sides of the equation, and add 1.9t to both sides of the equation:
18-18-3.8t=7036-1.9t-18 (Subtract 18)
-3.8t=7018-1.9t (Add 1.9t to both sides)
1.9t+ -3.8t=7018-1.9t+1.9t
This will leave you with the equation:
-1.9t=7018 (Divide each side of the equation by -1.9)
t=-70180/19
The answer to this equation is a fraction. T is equal to -70180/19